OpenStudy (anonymous):

Find the equation of the line passing through the point R(-3; 3) that is perpendicular to the line passing through the points P(-2;-2) and Q(1; 2).

OpenStudy (anonymous):

when the two lines are perpendicular, product of their slopes are equal to -1 in a ques u have been given two points P and Q of a line , from P and Q determine slope first after that u can determine the slope of the other line

OpenStudy (anonymous):

after finding the slope of required line u can determine the equation of a line using slope point form y-y1 = m (x-x1) where m = slope and (x1,y1) = points through which the line passses